A company just paid a dividend of $1.95 per share, and that dividend is expected to grow at a constant rate of 4.50% per year in the future. The company's beta is 1.65, the market risk premium is 8.5%, and the risk-free rate is 5.50%. What is the company's current stock price?

$12.72

$13.56

$14.53

$15.64

$16.94
Answer #1

First we need to find the required rate of return using the CAPM equation:

r_e = r_f + \beta \times (risk \ premium)

r_e = 5.50 + 1.65 \times (8.50)

r_e = 19.53 \%

A company just paid a dividend of $1.95 per share, and that dividend is expected to grow at a constant rate of 4.50% per year in the future.

Using the above rate, we can find the stock price:

P_0 = \frac{D_0 \times (1+g)}{r-g}

P_0 = \frac{1.95 \times (1+0.045)}{0.1953 - 0.045}

P_0 = \$ 13.56